### v1

@@ 1 Introduction
This paper focuse on the preconditioning of a monolithic scheme for the coupled Stokes-Darcy flow problem, which models the interaction between free fluid and porous media flows. It has been extensively studied in recent years due to its numerous applications, such as industrial filtration, water/gas management in fuel cells, and surface–subsurface interactions [9, 17]. The classical set of coupling conditions contain continuity of normal velocity and balance of normal and tangential forces, with the latter being described by the Beavers-Joseph-Saffman condition, see [3, 39].
@@
In this paper, we focus on the coupled, linear, stationary Stokes-Darcy model and the formulation where the coupling is enforced using Lagrange multiplier on the interface [27]. Different formulations of the Stokes-Darcy problem have been analyzed in the literature, e.g. the formulation with Darcy problem in the primal form [13] or fully mixed formulation [20]. The formulations have been targeted by discretization schemes, for example, finite volume method [18, 31, 40], (mixed) finite element method [10, 11, 37], and finite difference method [41, 22, 36]. Iterative methods for the resulting linear systems have been developed including domain decomposition approaches [14, 46, 4], multilevel methods [33, 29] or block preconditioners [5, 25, 44, 45].
@@
To our knowledge, an optimal monolithic preconditioner, designed using the operator preconditioning framework and robust with respect to both physical and discretization parameters, has not been accomplished for the Lagrange multiplier formulation of the coupled Stokes-Darcy problem. In this work, we rigorously establish the well-posedness of the problem, paying in particular special attention to the boundary conditions of the subproblems (cf. numerical experiments in [25]). As will become evident the coupled problem is not well-posed uniformly with respect to the parameters in certain boundary condition configurations. These cases are characterized by existence of a near-kernel space. In turn, in our analysis, we examine the inf-sup condition, with particular focus on how the inf-sup constant depends on the physical parameters, the entire solution space, and the complement of the near-kernel space.
@@
Using stability estimates from the continuous problem, we will analyze the convergence of the Minimal Residual (MINRES) method. It is well-known that the convergence of such a Krylov subspace iterative method is closely related to the spectral properties of the system matrix [2, 38, 34, 42]. In particular, achieving fast convergence requires clustered eigenvalues that are bounded away from zero. However, in near-kernel cases of the Stokes-Darcy problem, the MINRES method exhibits slow convergence or even stagnation, meaning the residual remains nearly unchanged over a substantial number of iterations, depending on the physical parameter values. To better understand this behavior, we analyze the convergence of the MINRES method in the presence of a “bad” or slow-to-converge eigenvalue of the preconditioned matrix. Building on this analysis, we employ the deflation technique [23, 43], which is commonly used to shift isolated clusters of small eigenvalues, resulting in a tighter eigenvalue distribution and a reduced condition number. In connection with the Stokes flow, [16] applied deflation preconditioners to speed up the slow convergence of standard block preconditioners when the problem domain is elongated, see also [15].
@@ 2 Preliminaries and Model
In this section, we first introduce the required notation, and then present the Stokes-Darcy model and derive weak formulation of the problem. Additionally, we outline the abstract framework proposed in [25] for analyzing the well-posedness of general coupled problems, emphasizing the key assumptions and functional settings required to ensure stability and solvability of the system.
@@ 2.1 Preliminaries
[…]
@@
For $\Gamma$ a subset of $\partial\Omega$, we define $H_{00}^{1/2}(\Gamma)$ as the space of all $w\in$ $H^{1/2}(\Gamma)$ whose extension by zero to $\partial\Omega$ belongs to $H^{1/2}(\partial\Omega)$. We also define $H_{00}^{-1/2}(\Gamma)$ to be the dual of $H_{00}^{1/2}(\Gamma)$. For the details on fractional and trace spaces, we refer to [19].
@@ 2.2 Model
(2.1) $\begin{cases}-\nabla\cdot\bm{\sigma}(\mathbf{u}_{S},p_{S})=\bm{f}_{S},\quad\nabla\cdot\mathbf{u}_{S}=0,&\text{ in }\Omega_{S},\\ \mathbf{u}_{S}=0,&\text{ on }\Gamma_{S}^{E},\\ -\bm{\sigma}(\mathbf{u}_{S},p_{S})\cdot\mathbf{n}_{S}=0,&\text{ on }\Gamma_{S}^{N}.\end{cases}$
@@
[…]
On the interface $\Gamma$, we have $\mathbf{n}_{S}=-\mathbf{n}_{D}$, and $\bm{\tau}=\mathbf{I}-(\mathbf{n}_{S}\otimes\mathbf{n}_{S})$ denotes a tangential projection onto $\Gamma$. We also define $\beta_{\tau}=\alpha_{\text{BJS}}\frac{\mu}{\sqrt{\bm{\tau}\cdot\bm{\kappa}\bm{\tau}}}$, where $\bm{\kappa}=\mu\mathbf{K}$ is the permeability, and $\alpha_{\text{BJS}}>0$ is the Beavers-Joseph-Saffman (BJS) slip coefficient. As for the physical meaning of the interface conditions, the first interface condition represents the mass conservation across $\Gamma$, the second one represents the balance of normal forces, and the third condition is the BJS condition [3, 39].
@@ 2.3 Weak formulation
where $\mathbf{H}_{0,E}^{1}(\Omega_{S}):=\left\{\mathbf{v}\in[H^{1}(\Omega_{S})]^{d}|\ \mathbf{v}=0\text{ on }\Gamma_{S}^{E}\right\}.$ The velocity-pressure spaces on $\Omega_{D}$ are
@@
where $\mathbf{H}_{0,E}(\operatorname{div};\Omega_{D}):=\left\{\mathbf{v}\in[L^{2}(\Omega_{D})]^{d}|\ \nabla\cdot\mathbf{v}\in L^{2}(\Omega_{D}),\ \mathbf{v}\cdot\mathbf{n}_{D}=0\text{ on }\Gamma_{D}^{E}\right\}$. In addition, let $\mathbf{V}=\mathbf{V}_{S}\times\mathbf{V}_{D}$ and $Q=Q_{S}\times Q_{D}$. For any $\mathbf{v}\in\mathbf{V},\,q\in Q$, we denote $\mathbf{v}_{i}=\mathbf{v}|_{\Omega_{i}},\,q_{i}=q|_{\Omega_{i}},\,i=S,D$. In particular, if we consider the coupled problem with pure Dirichlet/essential boundary conditions, that is, $\Gamma_{S}^{N}=\Gamma_{D}^{N}=\emptyset$, then the pressure space should be redefined as $Q/\mathbb{R}=\left\{q=(q_{S},q_{D})\in Q_{S}\times Q_{D}|\int_{\Omega}q\mathrm{~{}d}x=0\right\}.$ Following [25], the weighted norms of $\mathbf{V}$ and $Q$ are given by
@@
$\|\mathbf{v}_{S}\|_{\mathbf{V}_{S}}$ $:=\left(\mu\|\mathbf{v}_{S}\|_{1,S}^{2}+\beta_{\tau}\|\mathbf{v}_{S}\cdot\bm{\tau}\|_{\Gamma}^{2}\right)^{1/2},\quad\|\mathbf{v}_{D}\|_{\mathbf{V}_{D}}:=\left({K}^{-1}\|\mathbf{v}_{D}\|_{D}^{2}+{K}^{-1}\|\nabla\cdot\mathbf{v}_{D}\|_{D}^{2}\right)^{1/2}$
@@
(2.6a) $a(\mathbf{u},\mathbf{v})+b(\mathbf{v},(p,\lambda))$ $=(\bm{f}_{S},\mathbf{v}_{S}),\quad$ $\forall\,\mathbf{v}\in\mathbf{V},$
@@
$b(\mathbf{v},(p,\lambda))=-(p_{S},\nabla\cdot\mathbf{v}_{S})_{S}-(p_{D},\nabla\cdot\mathbf{v}_{D})_{D}+(\lambda,T_{n}\mathbf{v}_{S})_{\Gamma}-(\lambda,T_{n}\mathbf{v}_{D})_{\Gamma}.$
@@
To distinguish $T_{n}\mathbf{u}_{S}$ and $T_{n}\mathbf{u}_{D}$, we introduce $T_{S}\mathbf{u}_{S}=T_{n}\mathbf{u}_{S}$ and $T_{D}\mathbf{u}_{D}=-T_{n}\mathbf{u}_{D}$, which will be used in the following abstract framework. We also let $L^{2}_{t}(\Gamma)$ be the space of functions $\mathbf{u}$ on $\Omega$ such that $\mathbf{u}\cdot\bm{\tau}\in L^{2}(\Gamma)$. The weak form Eq. 2.6 induces the following linear operator
@@ 2.4 Abstract framework
According to the Brezzi theory [6], the well-posedness of the Eq. 2.8 requires the existence of constants $\alpha_{i},C_{i},\gamma_{i},D_{i}>0$ such that
@@
We remark that $T_{i}:\mathbf{V}_{i}\rightarrow\Lambda_{i}^{\prime}$ and hence $T_{S}\mathbf{u}_{S}+T_{D}\mathbf{u}_{D}\in\Lambda_{S}^{\prime}+\Lambda_{D}^{\prime}$. Therefore, $\lambda\in\Lambda_{S}\cap\Lambda_{D}$ since $\left(\Lambda_{S}\cap\Lambda_{D}\right)^{\prime}=\Lambda_{S}^{\prime}+\Lambda_{D}^{\prime}$. Following Theorem 3.1 in [25], we have the following theorem.
@@
Suppose the problems (2.8) satisfy the Brezzi conditions (2.9a)-(2.9d) in $\mathbf{V}_{i}\times Q_{i}\times\Lambda_{i},i=S,D$ and the coercivity condition (2.10). Then the coupled problem (2.12) is well posed in $\mathbf{W}=\mathbf{V}_{S}\times Q_{S}\times\mathbf{V}_{D}\times Q_{D}\times\left(\Lambda_{S}\cap\Lambda_{D}\right)$ in the sense that
@@
$\geq\alpha_{S}\left\|\mathbf{u}_{S}\right\|_{\mathbf{V}_{S}}^{2}+\alpha_{D}\left\|\mathbf{u}_{D}\right\|_{\mathbf{V}_{D}}^{2}$
@@
Finally, let us verify the inf-sup condition. Let $R_{Q_{i}}^{-1}:Q_{i}\rightarrow Q_{i}^{\prime},R_{\Lambda_{i}}^{-1}:\Lambda_{i}\rightarrow\Lambda_{i}^{\prime}$ be the inverse Riesz maps of their corresponding spaces. By the Riesz Representation Theorem, this is an isometry between $Q_{i}$ and $Q_{i}^{\prime}$, meaning that $(R_{Q_{i}}^{-1}q,q)=\|q\|_{Q_{i}}^{2}$. Given $\left(q_{S},q_{D},\phi\right)\in Q_{S}\times Q_{D}\times(\Lambda_{S}\cap\Lambda_{D})$, let $\mathbf{u}_{i}^{*},p_{i}^{*},\lambda_{i}^{*}$ be the solution of
@@ 3.1 Well-posedness of the Stokes-Darcy problem
Let $\Omega_{S}\subset\mathbb{R}^{d}$ be a bounded domain with boundary decomposition $\partial\Omega_{S}=\Gamma\cup\Gamma_{S}^{E}\cup\Gamma_{S}^{N}$, where the components are assumed to be of non-zero measure and $\Gamma_{S}^{E}\cap\Gamma=\emptyset$. We consider the Stokes problem (2.1) with the boundary conditions $\mathbf{u}_{S}|_{\Gamma_{S}^{E}}=0,\ \bm{\sigma}(\mathbf{u}_{S},p_{S})\cdot\mathbf{n}_{S}|_{\Gamma_{S}^{N}}=0$ and
[…]
@@
$a_{S}((\mathbf{u}_{S},p_{S},\lambda),(\mathbf{v}_{S},q_{S},\phi))$ $=2\mu(\bm{\epsilon}(\mathbf{u}_{S}),\bm{\epsilon}(\mathbf{v}_{S}))_{S}+\beta_{\tau}(T_{t}\mathbf{u}_{S},T_{t}\mathbf{v}_{S})_{\Gamma}-(p_{S},\nabla\cdot\mathbf{v}_{S})_{S}$
$\quad-(\nabla\cdot\mathbf{u}_{S},q_{S})_{S}+(\lambda,T_{S}\mathbf{v}_{S})_{\Gamma}+(T_{S}\mathbf{u}_{S},\phi)_{\Gamma},$
@@
In order to prove that the problem (3.4) is well-posed, we need to verify the conditions (2.9a)-(2.9d) and (2.10). The coercivity (2.9a), the continuity (2.9b) and (2.10) are promptly satisfied. The boundedness (2.9d) follows from the Cauchy-Schwarz inequality and trace inequality,
@@
(3.5) $\sup_{\mathbf{v}_{S}\in\mathbf{V}_{S}\backslash\{0\}}\frac{(B_{S}\mathbf{v}_{S},p_{S})+(T_{S}\mathbf{v}_{S},\lambda)}{\|\mathbf{v}_{S}\|_{\mathbf{V}_{S}}}\geq\gamma_{S}\left(\|p_{S}\|_{Q_{S}}^{2}+\|\lambda\|_{\Lambda_{S}}^{2}\right)^{1/2}.$
@@
[…]
@@
$a_{D}((\mathbf{u}_{D},p_{D},\lambda),(\mathbf{v}_{D},q_{D},\phi))$ $={K}^{-1}(\mathbf{u}_{D},\mathbf{v}_{D})_{D}-(p_{D},\nabla\cdot\mathbf{v}_{D})_{D}$
$\quad+(\nabla\cdot\mathbf{u}_{D},q_{D})_{D}+(\lambda,T_{D}\mathbf{v}_{D})_{\Gamma}+(T_{D}\mathbf{u}_{D},\phi)_{\Gamma},$
@@
In this lemma, we mainly establish the Darcy inf-sup condition on domain $\Omega_{D}$ for which we follow the same approach as in Lemma 3.1. The continuity and the coercivity conditions are obviously true. Let $\mathcal{D}_{\Gamma}$ be the Dirichlet-to-Neumann (DtN) map. For given $(p_{D},\lambda)$ in $Q_{D}\times\Lambda_{D}$, we construct $w\in H^{1}(\Omega_{D})$ as the solution of the following elliptic problem
@@
[…]
Integrating by parts for (3.11a), invoking Cauchy-Schwarz inequality, the property of the DtN map, the trace inequality, and Poincaré inequality, we have
@@
Then, the inf-sup condition Eq. 2.9c follows from (3.14) and (3.15). This completes the proof.
@@
which has a unique solution in $H^{1}(\Omega_{D})/\mathbb{R}$. Following the proof of Lemma 3.4, we can obtain well-posedness of the Darcy problem.
@@
(EE) $\Lambda=\frac{1}{\sqrt{\mu}}H_{00}^{-1/2}(\Gamma)\cap\sqrt{{K}}H^{1/2}(\Gamma),\,$ $\mathbf{W}=\mathbf{V}\times\Big{(}\frac{1}{\sqrt{\mu}}L^{2}\left(\Omega_{S}\right)\times\sqrt{{K}}L^{2}\left(\Omega_{D}\right)\times\Lambda\Big{)}/\mathbb{R},$

### v2

@@ 1 Introduction
This paper focuse on the preconditioning of a monolithic scheme for the coupled Stokes-Darcy flow problem, which models the interaction between free fluid and porous media flows. It has been extensively studied in recent years due to its numerous applications, such as industrial filtration, water/gas management in fuel cells, and surface–subsurface interactions [cao2010coupled, ervin2009coupled]. The classical set of coupling conditions contain continuity of normal velocity and balance of normal and tangential forces, with the latter being described by the Beavers-Joseph-Saffman condition, see [Beavers1967BoundaryCA, Saffman1971On].
@@
In this paper, we focus on the coupled, linear, stationary Stokes-Darcy model and the formulation where the coupling is enforced using Lagrange multiplier on the interface [layton2002coupling]. Different formulations of the Stokes-Darcy problem have been analyzed in the literature, e.g. the formulation with Darcy problem in the primal form [discacciati2004domain] or fully mixed formulation [gatica2011analysis]. The formulations have been targeted by discretization schemes, for example, finite volume method [fetzer2017conditions, masson2016coupling, schneider2020coupling], (mixed) finite element method [CAUCAO2017943, CaucaoGaticaOyarzúaŠebestová, rui2009unified], and finite difference method [shiue2018convergence, greif2023block, rui2020mac]. Iterative methods for the resulting linear systems have been developed including domain decomposition approaches [discacciati2007robin, vassilev2014domain, boon2020parameter], multilevel methods [mu2007two, luo2017uzawa] or block preconditioners [boon2022robust, Holter2020robust, Strohbeck2023Robust, Strohbeck2024Efficient]. In particular, Boon et al. [boon2022robust] developed robust monolithic solvers for the coupled primal Stokes–Darcy problem, assuming that the interface is in contact with portions of the external boundary where Neumann conditions are imposed on both subdomains. However, the authors also noted important differences in the norm of the Lagrange multiplier at the interface compared to the mixed-form Stokes–Darcy problem studied in [Holter2020robust], and extensive numerical tests were presented to support this observation.
@@
To our knowledge, an optimal monolithic preconditioner, designed using the operator preconditioning framework and robust with respect to both physical and discretization parameters, has not been accomplished for the Lagrange multiplier formulation of the coupled Stokes-Darcy problem. In this work, we rigorously establish the well-posedness of the problem, paying in particular special attention to the boundary conditions of the subproblems (cf. numerical experiments in [Holter2020robust]). As will become evident the coupled problem is not well-posed uniformly with respect to the parameters in certain boundary condition configurations. These cases are characterized by existence of a near-kernel space. In turn, in our analysis, we examine the inf-sup condition, with particular focus on how the inf-sup constant depends on the physical parameters, the entire solution space, and the complement of the near-kernel space.
@@
Using stability estimates from the continuous problem, we will analyze the convergence of the Minimal Residual (MINRES) method. It is well-known that the convergence of such a Krylov subspace iterative method is closely related to the spectral properties of the system matrix [barrett1994templates, saad2003iterative, nielsen2013analysis, simoncini2013superlinear]. In particular, achieving fast convergence requires clustered eigenvalues that are bounded away from zero. However, in near-kernel cases of the Stokes-Darcy problem, the MINRES method exhibits slow convergence or even stagnation, meaning the residual remains nearly unchanged over a substantial number of iterations, depending on the physical parameter values. To better understand this behavior, we analyze the convergence of the MINRES method in the presence of a “bad” or slow-to-converge eigenvalue of the preconditioned matrix. Building on this analysis, we employ the deflation technique [gutknecht2012spectral, soodhalter2020surveysubspacerecyclingiterative], which is commonly used to shift isolated clusters of small eigenvalues, resulting in a tighter eigenvalue distribution and a reduced condition number. In connection with the Stokes flow, [Dumitrasc_2024] applied deflation preconditioners to speed up the slow convergence of standard block preconditioners when the problem domain is elongated, see also [dobrowolski2003lbb].
@@ 2 Preliminaries and Model
In this section, we first introduce the required notation, and then present the Stokes-Darcy model and derive weak formulation of the problem. Additionally, we outline the abstract framework proposed in [Holter2020robust] for analyzing the well-posedness of general coupled problems, emphasizing the key assumptions and functional settings required to ensure stability and solvability of the system.
@@ 2.1 Preliminaries
For a domain $\Omega_{i}\subset\mathbb{R}^{d},d\in\{2,3\}$, the Sobolev space $H^{m}(\Omega_{i})=W^{2,m}(\Omega_{i})$ is defined in the usual way, with the norm $\|\cdot\|_{m,i}$ and seminorm $|\cdot|_{m,i}$, for any $i=\{S,D\}$. When $m=0$, we identify $H^{0}(\Omega_{i})=L^{2}(\Omega_{i})$ and use $\|\cdot\|_{i}$ for $\|\cdot\|_{0,i}$, with associated inner product denoted by $(\cdot,\cdot)_{i}$. For a general Hilbert space $V$, we denote its norm by $\lVert\cdot\rVert_{V}$. For the $L^{2}$ space, we use the simplified notation $\lVert\cdot\rVert$ for the norm and $(\cdot,\cdot)$ for the corresponding inner product. To emphasize the domain $\Omega$, we may explicitly write $(\cdot,\cdot)_{\Omega}$ and $\|\cdot\|_{\Omega}$. Finally, function spaces consisting of functions with mean value zero are represented as quotient spaces, e.g., $L^{2}(\Omega)/\mathbb{R}$ denotes the space of $L^{2}$ functions on $\Omega$ with zero mean.
@@
For $\Gamma$ a subset of $\partial\Omega$, we define $H_{00}^{1/2}(\Gamma)$ as the space of all $w\in$ $H^{1/2}(\Gamma)$ whose extension by zero to $\partial\Omega$ belongs to $H^{1/2}(\partial\Omega)$. We also define $H_{00}^{-1/2}(\Gamma)$ to be the dual of $H_{00}^{1/2}(\Gamma)$. For the details on fractional and trace spaces, we refer to [juan2007nonmathcing].
@@ 2.2 Model
(2.1) $\begin{cases}-\nabla\cdot\bm{\sigma}(\mathbf{u}_{S},p_{S})=\bm{f}_{S},\quad\nabla\cdot\mathbf{u}_{S}=0,&\text{ in }\Omega_{S},\\ \mathbf{u}_{S}=\mathbf{0},&\text{ on }\Gamma_{S}^{E},\\ -\bm{\sigma}(\mathbf{u}_{S},p_{S})\cdot\mathbf{n}_{S}=\mathbf{0},&\text{ on }\Gamma_{S}^{N}.\end{cases}$
@@
(2.3a) $\mathbf{u}_{S}\cdot\mathbf{n}_{S}+\mathbf{u}_{D}\cdot\mathbf{n}_{D}$ $=0,\quad$ $\text{ on }\Gamma,$
(2.3b) $-2\mu\bm{\epsilon}\left(\mathbf{u}_{S}\right)\mathbf{n}_{S}\cdot\mathbf{n}_{S}+p_{S}$ $=p_{D},\quad$ $\text{ on }\Gamma,$
(2.3c) $-2\mu\bm{\epsilon}\left(\mathbf{u}_{S}\right)\mathbf{n}_{S}\cdot\bm{\tau}$ $=\beta_{\tau}\mathbf{u}_{S}\cdot\bm{\tau},\quad$ $\text{ on }\Gamma.$
On the interface $\Gamma$, we have $\mathbf{n}_{S}=-\mathbf{n}_{D}$, and $\bm{\tau}=\mathbf{I}-(\mathbf{n}_{S}\otimes\mathbf{n}_{S})$ denotes a tangential projection onto $\Gamma$. We also define $\beta_{\tau}=\alpha_{\text{BJS}}\frac{\mu}{\sqrt{\bm{\tau}\cdot\bm{\kappa}\bm{\tau}}}$, where $\bm{\kappa}=\mu\mathbf{K}$ is the permeability, and $\alpha_{\text{BJS}}>0$ is the Beavers-Joseph-Saffman (BJS) slip coefficient. As for the physical meaning of the interface conditions, the first interface condition represents the mass conservation across $\Gamma$, the second one represents the balance of normal forces, and the third condition is the BJS condition [Beavers1967BoundaryCA, Saffman1971On].
@@ 2.3 Weak formulation
where $\mathbf{H}_{0,E}^{1}(\Omega_{S}):=\left\{\mathbf{v}\in[H^{1}(\Omega_{S})]^{d}\,|\ \mathbf{v}=0\text{ on }\Gamma_{S}^{E}\right\}$ and $\mathbf{L}^{2}_{t}(\Gamma)=\{\,\mathbf{u}\in[L^{2}(\Gamma)]^{d}\,|\ \mathbf{u}\cdot\bm{\tau}\in L^{2}(\Gamma)\,\}$. The velocity-pressure spaces on $\Omega_{D}$ are
@@
where $\mathbf{H}_{0,E}(\operatorname{div};\Omega_{D}):=\left\{\mathbf{v}\in[L^{2}(\Omega_{D})]^{d}|\ \nabla\cdot\mathbf{v}\in L^{2}(\Omega_{D}),\ \mathbf{v}\cdot\mathbf{n}_{D}=0\text{ on }\Gamma_{D}^{E}\right\}$. In addition, let $\mathbf{V}=\mathbf{V}_{S}\times\mathbf{V}_{D}$ and $Q=Q_{S}\times Q_{D}$. For any $\mathbf{v}\in\mathbf{V},\,q\in Q$, we denote $\mathbf{v}_{i}=\mathbf{v}|_{\Omega_{i}},\,q_{i}=q|_{\Omega_{i}},\,i=S,D$. In particular, if we consider the coupled problem with pure Dirichlet/essential boundary conditions, that is, $\Gamma_{S}^{N}=\Gamma_{D}^{N}=\emptyset$, then the pressure space should be redefined as $Q/\mathbb{R}=\left\{q=(q_{S},q_{D})\in Q_{S}\times Q_{D}|\int_{\Omega}q\mathrm{~d}x=0\right\}.$ Following [Holter2020robust], the weighted norms of $\mathbf{V}$ and $Q$ are given by
@@
$\|\mathbf{v}_{S}\|_{\mathbf{V}_{S}}$ $:=\left(\mu\|\mathbf{v}_{S}\|_{1,S}^{2}+\beta_{\tau}\|\mathbf{v}_{S}\cdot\bm{\tau}\|_{\Gamma}^{2}\right)^{1/2},\quad\|\mathbf{v}_{D}\|_{\mathbf{V}_{D}}:=\left({K}^{-1}\|\mathbf{v}_{D}\|_{D}^{2}+{K}^{-1}\|\nabla\cdot\mathbf{v}_{D}\|_{D}^{2}\right)^{1/2},$
@@
(2.6a) $a(\mathbf{u},\mathbf{v})-b(\mathbf{v},(p,\lambda))$ $=(\bm{f}_{S},\mathbf{v}_{S}),\quad$ $\forall\,\mathbf{v}\in\mathbf{V},$
@@
$b(\mathbf{v},(p,\lambda))=(\nabla\cdot\mathbf{v}_{S},p_{S})_{S}+(\nabla\cdot\mathbf{v}_{D},p_{D})_{D}-(T_{n}\mathbf{v}_{S},\lambda)_{\Gamma}+(T_{n}\mathbf{v}_{D},\lambda)_{\Gamma}.$
@@
To distinguish $T_{n}\mathbf{u}_{S}$ and $T_{n}\mathbf{u}_{D}$, we introduce $T_{S}\mathbf{u}_{S}=T_{n}\mathbf{u}_{S}$ and $T_{D}\mathbf{u}_{D}=-T_{n}\mathbf{u}_{D}$, which will be used in the following abstract framework. The weak form Eq. 2.6 induces the following linear operator
@@ 2.4 Abstract framework
According to the Brezzi theory [brezzi1974existence], the well-posedness of the Eq. 2.8 requires the existence of constants $\alpha_{i},C_{i},\gamma_{i},D_{i}>0$ such that
@@
We remark that $T_{i}:\mathbf{V}_{i}\rightarrow\Lambda_{i}^{\prime}$ and hence $T_{S}\mathbf{u}_{S}+T_{D}\mathbf{u}_{D}\in\Lambda_{S}^{\prime}+\Lambda_{D}^{\prime}$. Therefore, $\lambda\in\Lambda_{S}\cap\Lambda_{D}$ since $\left(\Lambda_{S}\cap\Lambda_{D}\right)^{\prime}=\Lambda_{S}^{\prime}+\Lambda_{D}^{\prime}$. Following Theorem 3.1 in [Holter2020robust], we have the following theorem.
@@
Suppose the problems (2.8) satisfy the Brezzi conditions (2.9b)-(2.9d) in $\mathbf{V}_{i}\times Q_{i}\times\Lambda_{i},i=S,D$ and the coercivity condition (2.10). Then the coupled problem (2.12) is well posed in $\mathbf{W}=\mathbf{V}_{S}\times Q_{S}\times\mathbf{V}_{D}\times Q_{D}\times\left(\Lambda_{S}\cap\Lambda_{D}\right)$ in the sense that
@@
Finally, let us verify the inf-sup condition (2.9c). Let $R_{Q_{i}}^{-1}:Q_{i}\rightarrow Q_{i}^{\prime},R_{\Lambda_{i}}^{-1}:\Lambda_{i}\rightarrow\Lambda_{i}^{\prime}$ be the inverse Riesz maps of their corresponding spaces. By the Riesz Representation Theorem, this is an isometry between $Q_{i}$ and $Q_{i}^{\prime}$, meaning that $(R_{Q_{i}}^{-1}q,q)=\|q\|_{Q_{i}}^{2}$. Given $\left(q_{S},q_{D},\phi\right)\in Q_{S}\times Q_{D}\times(\Lambda_{S}\cap\Lambda_{D})$, let $\mathbf{u}_{i}^{*},p_{i}^{*},\lambda_{i}^{*}$ be the solution of
@@ 3.1 Well-posedness of the Stokes-Darcy problem
Let $\Omega_{S}\subset\mathbb{R}^{d}$ be a bounded domain with boundary decomposition $\partial\Omega_{S}=\Gamma\cup\Gamma_{S}^{E}\cup\Gamma_{S}^{N}$, where the components are assumed to be of non-zero measure and $\Gamma_{S}^{E}\cap\Gamma=\emptyset$. We consider the Stokes problem (2.1) with the boundary conditions $\mathbf{u}_{S}|_{\Gamma_{S}^{E}}=\mathbf{0},\ \bm{\sigma}(\mathbf{u}_{S},p_{S})\cdot\mathbf{n}_{S}|_{\Gamma_{S}^{N}}=\mathbf{0}$ and
(3.1) $\mathbf{u}_{S}\cdot\mathbf{n}_{S}=g_{n},$  $\quad\text{ on }\Gamma,$
(3.2) $\bm{\tau}\cdot\left(2\mu\bm{\epsilon}(\mathbf{u}_{S})-p_{S}I\right)\cdot\mathbf{n}_{S}+{\beta_{\tau}}\mathbf{u}_{S}\cdot\bm{\tau}=0,$  $\quad\text{ on }\Gamma,$
@@
$a_{S}((\mathbf{u}_{S},p_{S},\lambda),(\mathbf{v}_{S},q_{S},\phi))$ $=2\mu(\bm{\epsilon}(\mathbf{u}_{S}),\bm{\epsilon}(\mathbf{v}_{S}))_{S}+\beta_{\tau}(T_{t}\mathbf{u}_{S},T_{t}\mathbf{v}_{S})_{\Gamma}-(\nabla\cdot\mathbf{v}_{S},p_{S})_{S}$
$\quad-(\nabla\cdot\mathbf{u}_{S},q_{S})_{S}+(T_{S}\mathbf{v}_{S},\lambda)_{\Gamma}+(T_{S}\mathbf{u}_{S},\phi)_{\Gamma},$
@@
In order to prove that the problem (3.4) is well-posed, we need to verify the conditions (2.9b)-(2.9d) and (2.10). The continuity (2.9b) and coercivity (2.10) are promptly satisfied. The boundedness (2.9d) follows from the Cauchy-Schwarz inequality and trace inequality,
@@
[…]
@@
(3.10) $\mathbf{u}_{D}\cdot\mathbf{n}_{D}=-g_{n},$  $\quad\text{ on }\Gamma,$
@@
$a_{D}((\mathbf{u}_{D},p_{D},\lambda),(\mathbf{v}_{D},q_{D},\phi))$ $={K}^{-1}(\mathbf{u}_{D},\mathbf{v}_{D})_{D}-(\nabla\cdot\mathbf{v}_{D},p_{D})_{D}$
$\quad+(\nabla\cdot\mathbf{u}_{D},q_{D})_{D}+(T_{D}\mathbf{v}_{D},\lambda)_{\Gamma}+(T_{D}\mathbf{u}_{D},\phi)_{\Gamma},$
@@
In this lemma, we mainly establish the Darcy inf-sup condition on domain $\Omega_{D}$ for which we follow the same approach as in Lemma 3.1. For given $(p_{D},\lambda)$ in $Q_{D}\times\Lambda_{D}$, we construct $w\in H^{1}(\Omega_{D})$ as the solution of the following elliptic problem
(3.11a) $-\nabla\cdot\nabla w={K}p_{D},$  $\quad\text{ in }\Omega_{D},$
(3.11b) $\nabla w\cdot\mathbf{n}_{D}={K}\mathcal{D}_{\Gamma}\lambda,$  $\quad\text{ on }\Gamma,$
(3.11c) $\nabla w\cdot\mathbf{n}_{D}=0,$  $\quad\text{ on }\Gamma_{D}^{E},$
(3.11d) $w=0,$  $\quad\text{ on }\Gamma_{D}^{N},$
where $\mathcal{D}_{\Gamma}:H^{1/2}_{00}(\Gamma)\rightarrow H^{-1/2}_{00}(\Gamma)$ is the Dirichlet-to-Neumann (DtN) mapping [sylvester1990dirichlet]. It is defined by $\mathcal{D}_{\Gamma}\lambda=\nabla u\cdot\mathbf{n}_{D}|_{\Gamma}$, where $u$ is the solution of
@@
Integrating by parts for (3.11a), invoking Cauchy-Schwarz inequality, the boundness-property of the DtN mapping, the trace inequality, and Poincaré inequality, we have
@@
Then, the inf-sup condition Eq. 2.9c follows from (3.15) and (3.16). Combining with he continuity and coercivity conditions completes the proof.
@@
which has a unique solution in $H^{1}(\Omega_{D})/\mathbb{R}$. Here, $\mathcal{D}_{\Gamma}:H^{1/2}(\Gamma)\rightarrow H^{-1/2}(\Gamma)$, the difference from (3.12) is that solution $u$ satisfies $\nabla u\cdot\mathbf{n}_{D}|_{\Gamma_{D}^{E}}=0$. Following the proof of Lemma 3.4, we can obtain well-posedness of the Darcy problem.
@@
($E$) $\Lambda=\frac{1}{\sqrt{\mu}}H_{00}^{-1/2}(\Gamma)\cap\sqrt{{K}}H^{1/2}(\Gamma),\,$ $\mathbf{W}=\mathbf{V}\times\Big(\frac{1}{\sqrt{\mu}}L^{2}\left(\Omega_{S}\right)\times\sqrt{{K}}L^{2}\left(\Omega_{D}\right)\times\Lambda\Big)/\mathbb{R},$